# EarthquakeNPP: A Benchmark for Earthquake Forecasting with Neural Point Processes

### v2

@@ A.2 Earthquake Catalog Completeness
All of the EarthquakeNPP datasets are made publicly available by their respective data centers in raw format. However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing. This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ [e.g., 42, 43].
There are several reasons why an earthquake may not be detected by a seismic network. Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations. Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high [34, 20]. Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g. when phase arrivals of different events overlap at different stations or the amplitudes of small events are swamped by those of large events. Since catalog incompleteness increases for lower magnitude events, typically the task is to find the value $M_{c}$ above which there is approximately $100\%$ detection probability. Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data. Where NPPs have demonstrated an ability to perform well with incomplete data [77], typically a threshold below the completeness biases classical models such as ETAS [67]. Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds [e.g. 40, 77], which we also facilitate in our datasets.
Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law [17], which states that the distribution of earthquake magnitudes follows an exponential probability density function
@@
where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$. Histogram-based approaches, such as the simple Maximum Curvature method [91] as well as many others [e.g. 26, and references therein], identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ [e.g. 65]. During aftershock sequences, $M_{c}(t)$ can be very high [e.g., 1, 19] (See Figure 6a). Thresholding at the maximum value might remove too much data. Instead, modelers either omit particularly incomplete periods during training and testing [32, 21], model the incompleteness itself [25, 89, 55, 18, 19, 45, 20], or accept known biases from disregarding this issue [73]. Spatially, catalogs are less complete farther from the seismic network [42], so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
@@ Appendix B Additional Datasets
Synthetic ETAS Catalogs. We simulate a synthetic catalog using the ETAS model with parameters estimated from ComCat, at $M_{c}\ 2.5$, within the same California region. A second catalog emulates the time-varying data-missingness present in observational catalogs by removing events using the time-dependent formula from Page et al. [59],
$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (9)
Deprecated Catalog of Japan. To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al. [5], however with a chronological train-test split and without removing any supposed outlier events. To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
@@
Appendix C Computational Efficiency
C.1 Training
@@
ETAS training scales $\mathcal{O}(n^{2})$ with the total number of events, since for every event a contribution to the intensity function is computed from a summation over all previous events. This scaling, coupled with the lack of parallelization in the current implementation, results in long training times for larger datasets. Poorer scaling will likely hinder ETAS if dataset sizes continue to grow in the future [76].
@@
Encouragingly, both DeepSTPP and AutoSTPP are significantly faster to train due to GPU acceleration and their use of a sliding window of the most recent $k=20$ events. While exact complexity analyses are not provided in Zhou et al. [97] or Zhou and Yu [96], we can infer that DeepSTPP likely scales as $\mathcal{O}(kn)$ since it benefits from a closed-form expression for the likelihood. AutoSTPP, though requiring automatic integration to compute the likelihood, still scales with $\mathcal{O}(kn)$ because the additional integration cost does not affect the overall scaling.
@@
NSTPP, on the other hand, incurs significant training costs, rendering it impractical for real-time forecasting. Unlike the sliding window mechanism used in DeepSTPP and AutoSTPP, NSTPP partitions the event sequence into fixed time intervals, leading to sequences that are much longer than the $k=20$ events used by the other models (as shown in Figure 11 of Chen et al. [5]). Furthermore, solving an ODE for each event time adds a significant computational burden, even with the use of their faster attentive CNF architecture.
@@
Whilst SMASH and DSTPP are built on the same backbone architecture, SMASH is much quicker to train than DSTPP—even faster than ETAS. This efficiency arises from its use of a single-step, normalization-free score-matching objective, which avoids the costly denoising and sampling loops required in diffusion-based training. SMASH directly learns the gradient of the log-density via pseudolikelihood estimation, enabling efficient GPU parallelization and bypassing the need for repeated evaluations over diffusion steps. In contrast, DSTPP simulates a sequential generative process over hundreds of intermediate steps per sample, significantly increasing computation and memory costs.
C.2 Simulation
Real-time earthquake forecasting and CSEP model evaluation require simulating many repeat sequences (at least 10,000 for adequate distributional coverage) over the forecasting horizon. While ETAS training scales as $\mathcal{O}(n^{2})$ with the number of training events, its simulation scales more efficiently at $\mathcal{O}(n\log n)$. This improved scaling is due to its equivalent formulation as a Hawkes branching process (see Section 2.2). Both Deep-STPP and AutoSTPP are also based on Hawkes processes, which theoretically allows for fast simulation. However, as these models currently only have an intensity function implementation, simulating events would require a slower thinning procedure [52], limiting their simulation efficiency. NSTPP too has slow simulation, since it must solve an ODE to sample a new event. Due to the slow simulation of DeepSTPP, AutoSTPP and NSTPP, they are not evaluated using the CSEP generative evaluation metrics.
Appendix D Analysis of Likelihood Scores
@@
(c) SanJac
(d) SaltonSea
(e) WHITE
Figure 10: Cumulative information gain (IG) plots for the spatial performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
Appendix E CSEP Consistency Tests
E.1 Number (Temporal) Test
@@
E.2 Pseudo-Likelihood Test
@@ E.4 Magnitude Test
E.5 Evaluating Multiple Forecasting Periods
Savran et al. [63] describe how to assess a model’s performance across the multiple days in the testing period (Figure 12). By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator [15]. Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast. Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively). The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
@@ E.5 Evaluating Multiple Forecasting Periods
Appendix F Further Dataset Figures
F.1 ComCat

### v3

@@ A.2 Earthquake Catalog Completeness
All of the EarthquakeNPP datasets are made publicly available by their respective data centers in raw format. However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing. This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ (e.g., Mignan et al., 2011; Mignan and Woessner, 2012).
There are several reasons why an earthquake may not be detected by a seismic network. Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations. Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high (Kagan and Knopoff, 1987; Hainzl, 2022). Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g. when phase arrivals of different events overlap at different stations or the amplitudes of small events are swamped by those of large events. Since catalog incompleteness increases for lower magnitude events, typically the task is to find the value $M_{c}$ above which there is approximately $100\%$ detection probability. Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data. Where NPPs have demonstrated an ability to perform well with incomplete data (Stockman et al., 2023), typically a threshold below the completeness biases classical models such as ETAS (Seif et al., 2017). Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds (e.g. Mancini et al., 2022; Stockman et al., 2023), which we also facilitate in our datasets.
Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law (Gutenberg and Richter, 1936), which states that the distribution of earthquake magnitudes follows an exponential probability density function
@@
where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$. Histogram-based approaches, such as the simple Maximum Curvature method (Wiemer and Wyss, 2000) as well as many others (e.g. Herrmann and Marzocchi, 2021, and references therein), identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ (e.g. Schorlemmer and Woessner, 2008). During aftershock sequences, $M_{c}(t)$ can be very high (e.g., Agnew, 2015; Hainzl, 2016b) (See Figure 6a). Thresholding at the maximum value might remove too much data. Instead, modelers either omit particularly incomplete periods during training and testing (Kagan, 1991; Hainzl et al., 2008), model the incompleteness itself (Helmstetter et al., 2006; Werner et al., 2011; Omi et al., 2014; Hainzl, 2016a; b; Mizrahi et al., 2021; Hainzl, 2022), or accept known biases from disregarding this issue (Sornette and Werner, 2005). Spatially, catalogs are less complete farther from the seismic network (Mignan et al., 2011), so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
@@ Appendix B Additional Datasets
B.1 Synthetic ETAS Catalogs.
We simulate a synthetic catalog using the ETAS model with parameters estimated from ComCat, at $M_{c}\ 2.5$, within the same California region. A second catalog emulates the time-varying data-missingness present in observational catalogs by removing events using the time-dependent formula from Page et al. (2016),
$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (9)
B.2 Deprecated Catalog of Japan.
To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al. (2021), however with a chronological train-test split and without removing any supposed outlier events. To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
We can use this corrected dataset to quantify the inflation of performance caused by the non-chronological training-validation-testing splitting in the Chen et al. (2021) dataset. Table 4 presents the information gain (difference in total log-likelihood, see section 2.1) relative to a Poisson process for the three NPP models across the two datasets. The dramatic drop in information gain highlights how the original data split and omission of the 2011 Tohoku earthquake inflates model performance.
Table 4: NPP performance comparison on the Original Chen et al. (2021) dataset versus the Japan_Deprecated dataset. Values are reported in terms of information gain from a homogeneous Poisson process.
Model Original Chen et al. (2021) dataset Japan_Deprecated
NSTPP 11.26 1.99
DeepSTPP 11.77 2.95
AutoSTPP 12.16 2.76
B.3 Likelihood Evaluation
@@
Appendix C Effect of Training Window on ETAS Performance
To verify that fitting ETAS on both the training and validation windows does not artificially improve its performance relative to the NPPs, we retrained ETAS using only the training window and re-evaluated its test log-likelihoods. As shown in Figures  9 & 10, the resulting log-likelihood scores are effectively unchanged across all EarthquakeNPP datasets.
Figure 9: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs. ETAS (orange) is trained on both training and validation windows, whereas ETAS* (light blue) is trained only using the training window.
Figure 10: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs. ETAS (orange) is trained on both training and validation windows, whereas ETAS* (light blue) is trained only using the training window.
Appendix D Computational Efficiency
D.1 Training
@@
ETAS training scales $\mathcal{O}(n^{2})$ with the total number of events, since for every event a contribution to the intensity function is computed from a summation over all previous events. This scaling, coupled with the lack of parallelization in the current implementation, results in long training times for larger datasets. Poorer scaling will likely hinder ETAS if dataset sizes continue to grow in the future (Stockman et al., 2024).
@@
Encouragingly, both DeepSTPP and AutoSTPP are significantly faster to train due to GPU acceleration and their use of a sliding window of the most recent $k=20$ events. While exact complexity analyses are not provided in Zhou et al. (2022) or Zhou and Yu (2024), we can infer that DeepSTPP likely scales as $\mathcal{O}(kn)$ since it benefits from a closed-form expression for the likelihood. AutoSTPP, though requiring automatic integration to compute the likelihood, still scales with $\mathcal{O}(kn)$ because the additional integration cost does not affect the overall scaling.
@@
NSTPP, on the other hand, incurs significant training costs, rendering it impractical for real-time forecasting. Unlike the sliding window mechanism used in DeepSTPP and AutoSTPP, NSTPP partitions the event sequence into fixed time intervals, leading to sequences that are much longer than the $k=20$ events used by the other models (as shown in Figure 11 of Chen et al. (2021)). Furthermore, solving an ODE for each event time adds a significant computational burden, even with the use of their faster attentive CNF architecture.
@@
Whilst SMASH and DSTPP are built on the same backbone architecture, SMASH is much quicker to train than DSTPP, even faster than ETAS. This efficiency arises from its use of a single-step, normalization-free score-matching objective, which avoids the costly denoising and sampling loops required in diffusion-based training. SMASH directly learns the gradient of the log-density via pseudolikelihood estimation, enabling efficient GPU parallelization and bypassing the need for repeated evaluations over diffusion steps. In contrast, DSTPP simulates a sequential generative process over hundreds of intermediate steps per sample, significantly increasing computation and memory costs.
D.2 Inference
Whilst log-likelihood computation for ETAS, Poisson, DeepSTPP, AutoSTPP, and NSTPP is fast (<30<30 seconds per dataset), real-time earthquake forecasting and CSEP evaluation require simulating many repeated event sequences (at least 10,000) over a fixed forecasting horizon. While ETAS training scales as $\mathcal{O}(n^{2})$ with the number of training events, its simulation is considerably more efficient, scaling approximately as $\mathcal{O}(n\log n)$ due to its equivalent formulation as a Hawkes branching process (see Section 2.2), which enables immigration–birth sampling.
Fast simulation is not currently feasible for several NPP architectures. NSTPP is not Hawkes-based and requires solving a neural ordinary differential equation to generate each new event, rendering simulation prohibitively slow even for small datasets. DeepSTPP and AutoSTPP, while inspired by Hawkes processes, employ non-stationary neural triggering kernels that depend on the full event history. As a result, standard immigration–birth sampling cannot be applied, since triggering relationships change after each event and new events cannot be generated independently or in parallel. Simulation via thinning is also problematic: AutoSTPP does not enforce monotonic or decaying kernels, meaning the conditional intensity $\lambda^{*}(t,\mathbf{x})$ cannot be safely upper-bounded, while DeepSTPP can in principle be simulated via thinning but is extremely slow in practice. For these reasons, DeepSTPP, AutoSTPP, and NSTPP are excluded from CSEP generative evaluation. For the models where large-scale simulation is tractable (ETAS, SMASH, and DSTPP), we report inference and simulation times in Table 6. Based on simulation times, DSTPP is fastest, followed by ETAS, with SMASH consistently the slowest across datasets, reflecting the fact that DSTPP samples events via a fixed-length diffusion process with closed-form updates, whereas SMASH relies on iterative Langevin dynamics requiring many gradient evaluations per event.
Table 6: Simulation time for a batch of 100 repeated simulations across datasets. Reported times correspond to individual forecast days and are summarised as minimum, median, mean, and maximum over the testing window, reflecting variation in daily event counts rather than runtime stochasticity. Times are formatted as $H$:$M$:$S$.
Dataset No. Days Daily Counts ETAS SMASH DSTPP
(min / med / mean / max) (min / med / mean / max) (min / med / mean / max) (min / med / mean / max)
ComCat 4764 0 / 3 / 4.59 / 1241 00:00:24 / 00:01:25 / 00:01:33 / 01:49:25 00:02:40 / 00:03:41 / 00:04:20 / 08:26:00 00:00:08 / 00:01:10 / 00:01:12 / 00:31:20
SaltonSea 731 0 / 2 / 5.61 / 292 00:00:19 / 00:01:20 / 00:01:39 / 00:26:38 00:02:15 / 00:03:17 / 00:04:45 / 02:00:57 00:00:06 / 00:01:08 / 00:01:14 / 00:08:12
SanJac 731 0 / 5 / 6.02 / 241 00:00:34 / 00:01:35 / 00:01:41 / 00:22:11 00:03:28 / 00:04:30 / 00:04:55 / 01:40:15 00:00:12 / 00:01:13 / 00:01:14 / 00:06:58
SCEDC 2191 0 / 3 / 5.96 / 2134 00:00:23 / 00:01:25 / 00:01:40 / 03:07:19 00:02:40 / 00:03:41 / 00:04:54 / 14:28:21 00:00:07 / 00:01:10 / 00:01:14 / 00:53:05
WHITE 1461 0 / 13 / 16.48 / 520 00:01:15 / 00:02:17 / 00:02:36 / 00:46:31 00:06:43 / 00:07:45 / 00:09:10 / 03:33:28 00:00:23 / 00:01:24 / 00:01:30 / 00:13:46
Appendix E Analysis of Likelihood Scores
E.1 Temporal Information Gain
To better interpret the temporal likelihood results presented in Figure 2, we analyse how model performance evolves over time in each dataset in Figure 11, presenting the cumulative information gain (log-likelihood difference) of each NPP model over ETAS.
In Figure 11(a) on the ComCat dataset, the largest decreases in relative performance occur during the two largest sequences in the testing window, namely the 2010 El Mayor Cucapah M7.2 and the 2019 Ridgecrest M7.1 earthquakes. ETAS performs strongly in these periods, likely because it incorporates magnitude scaling which enables it to model large aftershock cascades effectively. In contrast, the 2014 South Napa M6.0 and 2014 Offshore Eureka M6.8 do not produce such a marked drop in relative performance. This suggests that capturing the largest aftershock sequences is a key limitation of current NPP models, while during background periods NPPs perform relatively better due to their ability to capture non-stationarity, a property directly not modelled by ETAS. Despite the overall worse performance of NSTPP, the relative decrease during large events is not as sharp as the other models, suggesting superior performance.
A similar pattern appears in Figure 11(b) for the SCEDC dataset. All models show a sharp reduction in performance during the 2019 Ridgecrest M7.1 sequence. During the quieter period leading up to Ridgecrest, DeepSTPP performs comparably to ETAS. This again suggests that the lack of explicit magnitude conditioning limits NPP performance during large sequences and highlights NPP models ability to capture background non-staionarities.
Figures 11(c), 11(d) and 11(e) show results for the smaller regional datasets, SanJac, SaltonSea and White. These smaller magnitude, more regionally concentrated datasets, don’t display the same constrasting "background", "mainshock" behaviour present in ComCat and SCEDC. Although the overall performance of the NPP models is below that of ETAS, improvements occur during the 2017 Brawley swarm in the SaltonSea dataset, whereby DeepSTPP achieves higher temporal likelihood than ETAS. This behaviour aligns with the known difficulty ETAS has in modelling swarm-like sequences that are not initiated by a large mainshock.
@@ E.1 Temporal Information Gain
E.2 Spatial Information Gain
To better interpret the spatial likelihood results presented in Figure 3, we analyse how model performance evolves over space in each dataset in Figures 12, 13, visualising geographically where NPPs outperform or fall behind ETAS.
Figures 12 and 13 show the spatial distribution of log-likelihood information gain across all datasets. Across both figures, a consistent pattern emerges. ETAS performs best in regions dominated by large, magnitude driven mainshock–aftershock sequences, while NPP performance degrades in the immediate vicinity of these events. In contrast, NPPs tend to perform more competitively in regions characterised by spatially complex or diffuse seismicity.
In the larger regional catalogs shown in Figure 12, this distinction is most apparent in the ComCat dataset. Near the Ridgecrest and El Mayor–Cucapah sequences, NPPs exhibit reduced information gain relative to ETAS, consistent with the absence of explicit magnitude driven triggering. However, in the complex tectonic setting of the Mendocino Triple Junction, NPPs achieve comparatively strong spatial performance, with frequent positive information gain relative to ETAS. This region is characterised by interacting fault systems and persistent background activity rather than a single dominant mainshock, suggesting that NPPs are better suited to modelling such non-stationary and spatially heterogeneous seismic regimes.
Figure 13 presents results for the smaller regional datasets San Jacinto, Salton Sea, and White. These regions are dominated by lower magnitude seismicity and swarm like behaviour, and here NPPs again perform more competitively relative to ETAS. AutoSTPP assigns probability more smoothly along fault structures similarly to ETAS, resulting in information gain values concentrated near zero, while DeepSTPP and NSTPP show more heterogeneous behaviour. NSTPP produces more extreme and spikier information gain values that are spatially scattered, indicating less precise spatial localisation despite occasional high likelihood assignments. Instability during training led to the drastic underperformance of NSTPP on the White dataset and consequent low likelihood scores distributed across the entire region.
Overall, the spatial results mirror the temporal analysis. Current NPP architectures struggle most in mainshock dominated regimes but show clear promise in modelling spatially complex background seismicity and swarm driven activity, motivating future work on incorporating large magnitude triggering while preserving this flexibility.
@@ E.2 Spatial Information Gain
Figure 12: Spatial information gain per event, for NPP models relative to ETAS on (a) ComCat and (b) SCEDC. Scatter points correspond to the geographical location of the forecasted event, coloured by the value of the information gain over the ETAS model (green positive, red negative). For each model an inset plot displays the distribution of the spatial information gains for all events in the testing period.
(a) SanJac
(b) SaltonSea
(c) White
Figure 13: Spatial information gain per event, for NPP models relative to ETAS on (a) SanJac, (b) SaltonSea, and (c) White. Scatter points correspond to the geographical location of the forecasted event, coloured by the value of the information gain over the ETAS model (green positive, red negative). For each model an inset plot displays the distribution of the spatial information gains for all events in the testing period.
Appendix F CSEP Consistency Tests
F.1 Number (Temporal) Test
@@
F.2 Pseudo-Likelihood Test
@@ F.4 Magnitude Test
Although SMASH is, in principle, capable of modelling earthquake magnitudes, we restrict it to rate forecasting in this benchmark, as extending it to fine-grained magnitude prediction led to a deterioration in spatio-temporal performance metrics during training.
@@
F.5 Evaluating Multiple Forecasting Periods
Savran et al. (2020) describe how to assess a model’s performance across the multiple days in the testing period (Figure 15). By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator (Gneiting and Katzfuss, 2014). Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast. Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively). The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
@@ F.5 Evaluating Multiple Forecasting Periods
Appendix G Analysis of CSEP Tests
G.1 Temporal
To further interpret the CSEP consistency test results reported in Table 2, Figures 1618 show the daily event count forecasts produced by ETAS, SMASH, and DSTPP across all EarthquakeNPP datasets. These plots reveal systematic differences in how the models capture both background seismicity and sudden increases in earthquake rate.
Across all datasets, ETAS provides the most consistent forecasts of daily event counts. It captures low-activity days well and responds more effectively than the NPP-based models to increases in seismicity rate, although it still underpredicts the largest rate excursions associated with major earthquake sequences. This behaviour is particularly evident in the ComCat and SCEDC datasets during the 2010 El Mayor–Cucapah and 2019 Ridgecrest sequences (Figures 16(a) and 16(b)), and explains ETAS’s consistently high consistency test pass rates.
SMASH exhibits highly variable daily rate estimates across all regions. While this variability occasionally allows it to match periods of elevated seismicity, such as during El Mayor–Cucapah in ComCat, it more often leads to pronounced over and under prediction. This spiky behaviour results in a substantial number of failed consistency tests.
DSTPP produces much smoother daily rate forecasts than SMASH, but this comes at the cost of systematic underprediction. Across all datasets, DSTPP underestimates both background seismicity and elevated rate periods, with the effect becoming especially severe in SCEDC and White (Figures 16(b) and 18). This persistent bias explains its low consistency test pass rates.
Performance differences across datasets largely reflect the dominant seismic regime. In smaller regions such as San Jacinto and Salton Sea (Figures 17(a) and 17(b)), all models perform more competitively due to the absence of large mainshock-driven rate increases. However, even in these settings, SMASH remains overly variable and DSTPP continues to underestimate daily rates.
(a) ComCat
(b) SCEDC
Figure 16: Daily number forecasts from ETAS, SMASH, and DSTPP over the full testing period for (a) ComCat and (b) SCEDC. Vertical lines show the forecasted daily distributions of earthquake counts. Green lines indicate days where the observed count falls within the 95% forecast interval, while red lines indicate failures. Observed daily counts are shown as dots, with marker size proportional to the number of earthquakes and colour indicating pass (green) or fail (red).
(a) SanJac
(b) SaltonSea
Figure 17: Daily number forecasts from ETAS, SMASH, and DSTPP over the full testing period for (a) SanJac and (b) SaltonSea. Vertical lines show the forecasted daily distributions of earthquake counts. Green lines indicate days where the observed count falls within the 95% forecast interval, while red lines indicate failures. Observed daily counts are shown as dots, with marker size proportional to the number of earthquakes and colour indicating pass (green) or fail (red).
Figure 18: Daily number forecasts from ETAS, SMASH, and DSTPP over the full testing period for White. Vertical lines show the forecasted daily distributions of earthquake counts. Green lines indicate days where the observed count falls within the 95% forecast interval, while red lines indicate failures. Observed daily counts are shown as dots, with marker size proportional to the number of earthquakes and colour indicating pass (green) or fail (red).
G.2 Spatial
Figures 19 and 20 show aggregated spatial forecasts over the entire testing period for ETAS, SMASH, and DSTPP across all EarthquakeNPP datasets. These plots summarise how each model distributes seismicity rate spatially when forecasts are integrated over time. While this provides a useful overview of long-term spatial structure, it does not distinguish whether high rates are forecast before events occur.
Across all datasets, ETAS produces the most spatially concentrated forecasts, with high rates aligned along known fault structures and low rates assigned to seismically inactive regions. This apparent spatial precision arises largely from ETAS’s explicit modelling of clustering, whereby elevated rates are assigned in the vicinity of earthquakes that have already occurred. This behaviour is particularly evident in the ComCat and SCEDC datasets (Figures 19(a) and 19(b)), where ETAS reproduces the large-scale fault network and major clusters associated with past seismicity.
In contrast, SMASH consistently produces spatially diffuse forecasts. While it sometimes captures regions of elevated activity, its rate is spread broadly across each domain, leading to weaker spatial contrast between active and inactive regions. This effect is most pronounced in ComCat, where SMASH concentrates strongly around the southern end of the domain near the 2010 El Mayor–Cucapah sequence, while remaining diffuse elsewhere. Similar behaviour is observed in SCEDC, where SMASH is again dominated by the El Mayor region and assigns comparatively high rates across large areas of the domain.
DSTPP generally produces smoother spatial forecasts than ETAS but with greater structure than SMASH. In ComCat, San Jacinto, and Salton Sea (Figures 19(a), 20(a), and 20(b)), DSTPP follows the main fault-aligned clustering while attenuating sharp spatial contrasts. However, in SCEDC and White, DSTPP substantially underestimates the overall seismicity rate, resulting in uniformly low spatial forecasts and visible boundary effects, particularly near the edges of the domains (Figures 19(b) and 20(c)).
(a) ComCat
(b) SCEDC
Figure 19: Spatial forecasts from ETAS, SMASH and DSTPP, aggregated across the entire testing period of (a) ComCat and (b) SCEDC. For each day in the testing period, every model simulates 10,000 repeated earthquake catalogs within the boundary region. All simulated catalogs are aggregated across the entire testing period and the expected earthquake rates are plotted within each grid cell. Observed earthquakes are overlaid with marker size proportional to magnitude.
(a) SanJac
(b) SaltonSea
(c) White
Figure 20: Spatial forecasts from ETAS, SMASH and DSTPP, aggregated across the entire testing period of (a) SanJac, (b) SaltonSea, and (c) White. For each day in the testing period, every model simulates 10,000 repeated earthquake catalogs within the boundary region. All simulated catalogs are aggregated across the entire testing period and the expected earthquake rates are plotted within each grid cell. Observed earthquakes are overlaid with marker size proportional to magnitude.
Appendix H Further Dataset Figures
H.1 ComCat